OpenStudy (anonymous):

Solve: 3(x + 1) + 6 = 33 10 8 8.7 −8

OpenStudy (zepp):

You need to solve x, so let's isolate it \(3(x+1)=33-6\\3(x+1)=27\\(x+1)=27\div3\\(x+1)=9\\x=9-1\\x=8\) Do you understand these steps?

OpenStudy (anonymous):

Yes!

OpenStudy (zepp):

Good :)

OpenStudy (anonymous):

Thanks So Much!

OpenStudy (zepp):

You are welcome!